# Blockchain Data Analytics: Review and Challenges

### v1

@@ 4.3 Blockchain Research Platforms
In addition to academic research, industry-driven blockchain intelligence platforms provide essential analytics for a broad range of users, including individual investors, institutional clients, regulatory bodies, and financial analysts. Messari [62] offers proprietary analytics dashboards that track blockchain projects, DeFi protocols, and token performance, supplemented by AI-powered sentiment analysis for assessing market trends and investor sentiment. Additionally, it provides institutional-grade research reports on emerging blockchain technologies, governance models, and regulatory developments. Nansen [66] specializes in tracking smart money flows, offering insights into the movements of institutional and high-net-worth investors, as well as on-chain behavior patterns of individual traders. Chainalysis [16] is a leading platform for blockchain forensic analysis, focusing on fraud detection, anti-money laundering (AML) compliance, and tracking illicit financial activities across blockchain networks. Token Terminal [78] provides fundamental financial data on crypto projects, analyzing revenue generation, protocol fees, and user adoption metrics. IntoTheBlock and Glassnode offer blockchain-based trading analytics, equipping investors with on-chain indicators, trading signals, and predictive models to optimize investment strategies. Bubblemaps [11] utilizes a graph-based visualization approach to represent token holder distributions, where wallets are displayed as interconnected bubbles, allowing users to track ownership concentration and potential market manipulation. These platforms collectively enhance blockchain intelligence by providing valuable tools for market research, financial analysis, and regulatory compliance.
@@ 5.1 Accessibility
Another significant accessibility challenge is the raw and unstructured nature of blockchain data. Public blockchain networks such as Bitcoin and Ethereum store transactional data in low-level formats, requiring additional processing before meaningful analysis can be conducted. For instance, smart contract interactions on Ethereum must be decoded using the Contract Application Binary Interface (ABI) to retrieve structured data. To simplify this process, indexing services such as The Graph [35] provide structured APIs that enable efficient data retrieval, eliminating the need for direct decoding. However, these platforms primarily focus on specific ecosystems, such as EVM-compatible chains, which limits their applicability to non-EVM blockchains. Other platforms, such as Dune Analytics [29], offer broader multi-chain support and allow users to query blockchain data using SQL. However, executing complex SQL queries can be time-consuming, and in some cases, users may need to upgrade to premium services to improve query performance. An alternative approach is to access blockchain datasets available through Google BigQuery [^3] https://cloud.google.com/blockchain-analytics/docs/supported-datasets, which provides structured blockchain data for analysis using SQL-based queries. For users who prefer self-hosted solutions, blockchain data can be processed using open-source ETL (Extract, Transform, Load) frameworks such as the Blockchain ETL project [^4] https://github.com/blockchain-etl, which facilitates structured extraction from raw blockchain data.
Overall, blockchain data accessibility has improved significantly, with services ranging from raw data providers to advanced analytics platforms such as Messari [62] and Nansen [66], which offer ready-to-use insights. However, the most reliable and high-performance data services often require a subscription, limiting access for researchers with constrained resources. Another key challenge is enhancing the interpretability of blockchain data, particularly for beginners. Advanced data visualization techniques, such as those proposed by SilkViser [91], can improve the presentation of blockchain data, making it more intuitive for users with varying levels of expertise. Additionally, automated visualization recommendation systems, similar to the approaches introduced by Mafrur et al. [58, 75, 59], can facilitate the extraction of meaningful insights by dynamically adapting to user needs. Future research should focus on developing more effective visualization methods to enhance blockchain data accessibility for beginners, intermediate users, and advanced analysts alike.
@@ 5.3 Accuracy
As shown in Table 1, both academic research and industry blockchain analytics tools offer a wide range of solutions, from processing raw blockchain data to executing SQL queries or simply consuming pre-analyzed insights. However, a fundamental question arises: how can we verify the accuracy of these outputs? Are the insights generated by these platforms reliable, and who is responsible for their validation? Blockchain data is often subject to inconsistencies due to issues such as data duplication, erroneous smart contracts, and incomplete transaction records. Additionally, user-generated queries on open analytics platforms may introduce biases or inaccuracies, impacting the reliability of insights. As highlighted by [60], data quality plays a crucial role in visual analytics, directly affecting the accuracy and interpretability of analytical results. Ensuring the integrity of blockchain data analytics is essential for deriving meaningful insights and supporting informed decision-making.

### v2

@@ 4.3 Blockchain Research Platforms
In addition to academic research, industry-driven blockchain intelligence platforms provide essential analytics for a broad range of users, including individual investors, institutional clients, regulatory bodies, and financial analysts. Messari [63] offers proprietary analytics dashboards that track blockchain projects, DeFi protocols, and token performance, supplemented by AI-powered sentiment analysis for assessing market trends and investor sentiment. Additionally, it provides institutional-grade research reports on emerging blockchain technologies, governance models, and regulatory developments. Nansen [67] specializes in tracking smart money flows, offering insights into the movements of institutional and high-net-worth investors, as well as on-chain behavior patterns of individual traders. Chainalysis [16] is a leading platform for blockchain forensic analysis, focusing on fraud detection, anti-money laundering (AML) compliance, and tracking illicit financial activities across blockchain networks. Token Terminal [79] provides fundamental financial data on crypto projects, analyzing revenue generation, protocol fees, and user adoption metrics. IntoTheBlock and Glassnode offer blockchain-based trading analytics, equipping investors with on-chain indicators, trading signals, and predictive models to optimize investment strategies. Bubblemaps [11] utilizes a graph-based visualization approach to represent token holder distributions, where wallets are displayed as interconnected bubbles, allowing users to track ownership concentration and potential market manipulation as shown in Figure 1. These platforms collectively enhance blockchain intelligence by providing valuable tools for market research, financial analysis, and regulatory compliance.
Figure 1: Bubblemap visualization
@@ 5.1 Accessibility
A significant accessibility challenge lies in the raw and unstructured nature of blockchain data. Public blockchain networks such as Bitcoin and Ethereum store transactional data in low-level formats, requiring additional processing before meaningful analysis can be conducted. For instance, smart contract interactions on Ethereum must be decoded using the Contract Application Binary Interface (ABI) to retrieve structured data. To simplify this process, indexing services such as The Graph [35] provide structured APIs that enable efficient data retrieval, eliminating the need for direct decoding. However, these platforms primarily focus on specific ecosystems, such as EVM-compatible chains, which limits their applicability to non-EVM blockchains. Other platforms, such as Dune Analytics [29], offer broader multi-chain support and allow users to query blockchain data using SQL. However, executing complex SQL queries can be time-consuming, and in some cases, users may need to upgrade to premium services to improve query performance. An alternative approach is to access blockchain datasets available through Google BigQuery [^3] https://cloud.google.com/blockchain-analytics/docs/supported-datasets, which provides structured blockchain data for analysis using SQL-based queries. For users who prefer self-hosted solutions, blockchain data can be processed using open-source ETL (Extract, Transform, Load) frameworks such as the Blockchain ETL project [^4] https://github.com/blockchain-etl, which facilitates structured extraction from raw blockchain data.
Overall, blockchain data accessibility has improved significantly, with services ranging from raw data providers (e.g., QuickNode [73], Alchemy [1], Infura [43], and Chainstack [17]) to advanced analytics platforms such as Messari [63] and Nansen [67], which offer ready-to-use insights. However, the most reliable and high-performance data services often require a subscription, limiting access for researchers with constrained resources. Another key challenge is enhancing the interpretability of blockchain data, particularly for beginners. Advanced data visualization techniques, such as those proposed by SilkViser [92], can improve the presentation of blockchain data, making it more intuitive for users with varying levels of expertise. Additionally, automated visualization recommendation systems, such as those proposed by Mafrur et al. [58, 76, 59], can automatically extract and recommend meaningful insights from blockchain data. Future research should focus on developing more effective visualization methods to enhance blockchain data accessibility for beginners, intermediate users, and advanced analysts alike.
@@ 5.3 Accuracy
As shown in Table 1, both academic research and industry blockchain analytics tools offer a wide range of solutions, from processing raw blockchain data to executing SQL queries or simply consuming pre-analyzed insights. However, a fundamental question arises: how can we verify the accuracy of these outputs? Are the insights generated by these platforms reliable, and who is responsible for their validation? Blockchain data is often subject to inconsistencies due to issues such as data duplication, erroneous smart contracts, and incomplete transaction records. Additionally, user-generated queries on open analytics platforms may introduce biases or inaccuracies, impacting the reliability of insights. As highlighted by  [60, 61], data quality plays a crucial role in visual analytics, directly affecting the accuracy and interpretability of analytical results. Ensuring the integrity of blockchain data analytics is essential for deriving meaningful insights and supporting informed decision-making.